# An Incremental Unified Framework for Small Defect Inspection

### 14 Dec 2023 v1

@@ 5.2 Experiment Performance
Qualitative Evaluation Table 1 (A) (B) shows that our method achieves SOTA performance with different experiment settings in pixel-level (A) and Image-level (B), respectively. We observe that the network performance is catastrophic forgetting in the reconstruction-based approach (UniAD [34]). While some continuous learning strategies [36, 16, 2, 31] can partially mitigate this problem, our approach provides an optimal solution. Besides, compared with CAD-based methods, our method can offer not only pixel-level location but also SOTA performance for incrementing more objects. Overall, our algorithm can overcome the above drawbacks and maintain a low level of forgetting.
@@
Pixel-level AUROC (%) Image-level AUROC (%)
$\mathbf{14-1\ (1\ Step)}$ $\mathbf{10-5\ (1\ Step)}$ $\mathbf{3\times 5\ (5\ Steps)}$ $\mathbf{10-1\times 5\ (5\ Steps)}$ $\mathbf{14-1\ (1\ Step)}$ $\mathbf{10-5\ (1\ Step)}$ $\mathbf{3\times 5\ (5\ Steps)}$ $\mathbf{10-1\times 5\ (5\ Steps)}$
ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$)
w/o OASA 92.5 8.2 90.6 13.5 88.9 10.8 90.3 8.6 89.3 13.3 87.6 16.5 81.4 12.2 84.2 14.2
w/o SCL 96.1 0.8 94.3 6.0 88.3 11.7 95.1 2.7 95.5 2.4 92.0 8.1 82.7 10.6 93.4 3.6
w/o US 96.2 0.7 93.3 8.5 89.5 10.3 90.5 8.4 95.9 1.1 88.6 16.5 83.5 10.9 85.6 13.5
Ours 96.3 0.6 94.4 6.3 91.1 8.4 95.1 1.0 96.0 1.0 92.2 9.3 84.2 10.0 94.2 3.2
Table 2: Qualitative evaluation in ablation study. Red represents the best ACC. The results show that all components in our method, Object-Aware Self-Attention, semantic compression loss, and our updating strategy, contribute to performance improvement of the whole framework.
Figure 7: Qualitative evaluation of our updating strategy. Here are two location heatmaps in old objects. Compared with “w/o US” (In Red Box), it is obvious that our updating strategy (In Green Box) can be more accurate to locate defects.
@@ 5.3 Ablation Study
Effectiveness of Object-Aware Self-Attention To demonstrate the effectiveness of classifying attributes, we ablate Object-Aware Self-Attention in our structure. “w/o OASA” in Table 2 shows that this structure can reduce catastrophic forgetting by classifying information insertion.
Effectiveness of Semantic Compression Loss To demonstrate the effectiveness of Semantic Compression Loss, we ablate this loss in our training. “w/o SCL” in Table 2 shows that this structure can reduce catastrophic forgetting by preserving more space for new objects.
Effectiveness of Our Updating Strategy To demonstrate the effectiveness of our updating strategy, we ablate this updating method, and “w/o US” in Table 2 shows that this method can reduce catastrophic forgetting. Also, in Fig. 7, our algorithm can help memorize important semantic spaces for old objects, thus reducing the rewriting of semantic space in the reconstruction network.
@@ 6 Limitations
Firstly, our approach does not explicitly represent the semantic space capacity, which is unfavorable for measuring the network ability for object-incremental learning. Besides, feature sharing is also significant in incremental learning, feature sharing on the reconstruction network can reduce the space occupation of individual objects. These limitations are valuable to be further discussed.

### 24 Jan 2024 v2

@@ 5.2 Experiment Performance
Quantitative Evaluation Table 1 (A) (B) shows that our method achieves SOTA performance with different experiment settings in pixel-level (A) and Image-level (B), respectively. We observe that the network performance is catastrophic forgetting in the reconstruction-based approach (UniAD [34]). While some continuous learning strategies [36, 16, 2, 31] can partially mitigate this problem, our approach provides an optimal solution. Besides, compared with CAD-based methods, our method can offer not only pixel-level location but also SOTA performance for incrementing more objects. Overall, our algorithm can overcome the above drawbacks and maintain a low level of forgetting.
@@ 5.3 Ablation Study
[…]
ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$) ACC($\uparrow$) FM($\downarrow$)
w/o OASA 89.3 / 92.5 13.3 / 8.2 87.6 / 90.6 16.5 / 13.5 81.4 / 88.9 12.2 / 10.8 84.2 / 90.3 14.2 / 8.6
w/o SCL 95.5 / 96.1 2.4 / 0.8 92.0 / 94.3 8.1 / 6.0 82.7 / 88.3 10.6 / 11.7 93.4 / 95.1 3.6 / 2.7
w/o US 95.9 / 96.2 1.1 / 0.7 88.6 / 93.3 16.5 / 8.5 83.5 / 89.5 10.9 / 10.3 85.6 / 90.5 13.5 / 8.4
Ours 96.0 / 96.3 1.0 / 0.6 92.2 / 94.4 9.3 / 6.3 84.2 / 91.1 10.0 / 8.4 94.2 / 95.1 3.2 / 1.0
Table 2: Quantitative evaluation in ablation study. The results show that all components in our method, Object-Aware Self-Attention, semantic compression loss, and our updating strategy, contribute to performance improvement of the whole framework. Red and Gray Red represents the best image-level and pixel-level performance, respectively.
We conduct the ablation study on MvTec AD [3] to evaluate the effectiveness of different components in our proposed method.
Effectiveness of Object-Aware Self-Attention To demonstrate the effectiveness of classifying attributes, we ablate Object-Aware Self-Attention in our structure. “w/o OASA” in Table 2 shows that this structure can reduce catastrophic forgetting by classifying information insertion. Moreover, in Fig. 7 (“w/o OASA”), if there is no semantic boundary, it will lead to serious feature conflict, and then impact the performance of the reconstruction network.
Effectiveness of Semantic Compression Loss To demonstrate the effectiveness of Semantic Compression Loss, we ablate this loss in our training. “w/o SCL” in Table 2 shows that this structure can reduce catastrophic forgetting by preserving more space for new objects. Besides, in Fig. 7 (“w/o SCL”), by compressing the space, we obtain clearer defect localization, which indicates that more network space can be obtained for incremental learning.
Effectiveness of Our Updating Strategy To demonstrate the effectiveness of our updating strategy, we ablate this updating method, and “w/o US” in Table 2 shows that this method can reduce catastrophic forgetting. Also, in Fig. 7 (“w/o US”), our algorithm can help memorize important semantic spaces for old objects, thus reducing the rewriting of semantic space in the reconstruction network.
Input w/o OSOA w/o US w/o SCL Ours Ground Truth
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Qualitative evaluation of ablation study. Compared with “w/o US” and “w/o OSOA”, it is obvious that our method can be more accurate to locate defects. Compared with “w/o SCL”, since increasing the available space of the model, our algorithm can further reduce the feature conflict and reduce more interference in the background (Please zoom in for more details).
@@ 6 Limitations
Firstly, our approach does not explicitly represent the semantic space capacity, which is unfavorable for measuring the network ability for object-incremental learning. Besides, feature sharing is also significant in incremental learning since feature sharing on the reconstruction network can reduce the space occupation of individual objects. These limitations are valuable and should be further discussed.